RC Update and Run Plan for Aug 6, 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

Thursday Aug 6 2020.

  • Halls A/B are doing the pass change.
  • Hall C in beam permit waiting to re-establish beam.

General updates

  • Re-established beam around mid-owl shift. Max reliable current ~10 uA.
  • Accomplished Simona's SoLID Cerenkov test at 1,1.5,2 uA. [1]
  • Tried to run on 3He but could not due to target motion FSD [2], [3]
  • Ran the rest of shift and half of day shift on reference cell, all at ~10 uA
  • HMS dipole tripped due to CHL [4]. We continued running on SHMS during HMS recovery.
  • Ethan fixed the 3He FSD problem by reverting the target ladder position. [5] Target is now 1mm lower.
  • Observed ramp rates that were not consistent with the software. Experts checked and verified software is doing the right thing, but EPICs only refreshing every 3s so some jumps seem more uneven. [6]

Pending problem

  • Need higher current beam.
  • Alarming coil current limits: [7],[8]

Recent polarization measurements

Current Run Plan

Normal Production running

  • Prerequisites
    • Make sure the raster size is correct (4-4.5)
    • Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
    • Running on cell targets: raster always ON (4-4.5, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
    • Always take 120Hz of EDTM
    • Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
    • Change the IHWP every 4 hours when running on glass cells
  • Run plan
    • Re-establish beam [10].
    • If MCC can deliver up to 42 muA we will do a BCM calibration [11]
    • Production on 3He; if not able to reach 30 uA, pro-rate the ABUs

Notes for tomorrow's 7:45 meeting

  •  ?


  • Planning for Simona's SoLID Cerenkov access on Tuesday, will need to coordinate to rotate HMS to 29 deg, crane work, etc. ~4-5 hours work
  • Target polarization calibration and studies on Tuesday-Wednesday during beam off and Moller measurement
  • Moller measurement on Wednesday
  • After ~ 10 days of running Hall C will go to 1 pass for ~3 days

Hall Access

  • None planned right now

When there is a down:

  • If the down is 2+ hours call Simona: she will check with the RC to see if there is an opportunity for work in the Hall (SoLID Cherenkov parasitic test)

Notices in Effect for Shift Crew

  • Beam positions: [12]
  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[13]

Planned Accelerator Configuration Changes

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links